# Non-absolute integration and application to Young geometric integration

### v1

@@ 1. Introduction
The classical Young integral [46] is recovered when $d=1$. This integral has recently attracted noticeable attention. For instance, [1] presents an alternative construction of the 2-dimensional Züst integral. Most research, starting from Züst’s original work, has focused on this construction because of its connections with the Gromov-Hölder problem for the Heisenberg group; see, [25, 3].
@@ 2.3. Henstock integral in higher dimension
A natural question is whether one can formulate a divergence theorem for the Henstock integral that is valid for every differentiable vector field. Such a result would constitute a higher-dimensional analog of the fundamental theorem of calculus. Unfortunately, the answer is negative. This non-absolute integral does, however, satisfy a Fubini theorem (see [35, Section 6.1], and a remarkable argument due to W. Pfeffer shows that a gauge integral cannot simultaneously satisfy both a generalized Fubini theorem and divergence theorem, see [38, Example 11.1.2].
@@ 2.4. Charges
The first appearance of charges in the study of non-absolute integrals can be traced back to Mařík [33]. The condition $\sup_{n}\|B_{n}\|<\infty$ in the continuity statement ensures that the $B_{n}$ do not fractalize excessively. It is inspired by continuity properties of the Henstock-Kurzweil integral (see for instance Hake’s theorem [38, Theorem 6.1.6]). Note also that it implies that $\omega(B)=0$ whenever $|B|=0$.
@@ 2.6. Pfeffer integral
$\left|\mathscr{S}(f,\mathscr{P})-g(1)\right|\leq\sum_{i=0}^{n}\left|f(x_{i})(b_{i}-a_{i})-(g(b_{i})-g(a_{i}))\right|\leq\varepsilon.$
@@
The formulation is of Saks-Henstock type because, in general, $\delta$-fine partitions do not always exist. For example, when $d=1$ and $\delta(x)=x/2$ on $[0,1]$, it is impossible to construct a $\delta$-fine partition. However, the well-definedness of this definition is based on a powerful Cousin lemma, due to Howard ([29, Lemma 5] or [40, Lemma 2.6.4]), that guarantees the existence of almost filling $\delta$-fine $\eta$-regular partitions (for $\eta$ smaller than the regularity of cubes), and therefore, the uniqueness of the indefinite integral $\omega$. It is also possible to give a formulation of the Pfeffer integral in terms of Riemann sums, with control over the size of the residual set.
@@
In one dimension, the Pfeffer integral is strictly weaker than the Henstock-Kurzweil integral, essentially because compact intervals are replaced by finite unions of compact intervals, [37, Example 6.9]. This is a drawback, in view of the well-established status of the Henstock-Kurzweil integral. In dimension $d\geq 2$, the Henstock and Pfeffer integrals give rise to different classes of integrable functions, with no inclusion relation between them.
@@ 3.3. A multidimensional Faber-Schauder basis
as in Theorem 3.5. It is possible to make sense of the coefficients $\omega(h_{-1})$ and $\omega(h_{n,k,e})$, because the Haar functions are piecewise constant on dyadic cubes. Moreover, the Hölder control yields estimate of the coefficients $|\omega(h_{n,k,e})|$. Using (3.6), we then show that the series in (3.9) converges in $CH([0,1]^{d})$. Consequently, the left-hand side of (3.9) is well-defined as an element of $CH([0,1]^{d})$ and provides the desired extension. Proving that $\omega$ is actually $\gamma$-fractional requires an interpolation argument that is to be found in [9, Theorem 3]. ∎
@@
Therefore, the space $CH^{\gamma}([0,1]^{d})$ is isomorphic to $\ell^{\infty}$, in the same way that $C^{\gamma}([0,1])$ is related to $\ell^{\infty}$ via the Ciesielski isomorphism. In particular, $CH^{\gamma}([0,1]^{d})$ is a dual Banach space (indeed, a bidual, see [8, Proposition 7.2]).
@@
Finally, we briefly discuss the characterization of $\gamma$-fractional charges. It was shown in [9, Theorem 11] that they are precisely the divergences of $\gamma$-Hölder continuous vector fields. More specifically, the operator $\operatorname{div}$ defined in (3.1) induces a surjective map
@@ 4.1. Whitney’s theory of flat chains and flat cochains
This turns out to be the right choice, as (1) and (2) are easily established. The only difficult step is to prove $\mathbf{F}(T)=0\implies T=0$, which is carried out in [45, Chapter V, §12]. The norm $\mathbf{F}$ is called the flat norm. The following example shows that the flat norm encodes some sort of weak convergence, and why axiom (1) is essential.
@@ 4.3. Fractional currents and fractional charges
Comparing with (4.4), we guess that $\mathbf{CH}^{m,\gamma}([0,1]^{d})$ is the dual of $\mathbf{F}_{m}^{\gamma}([0,1]^{d})$, [8, Theorem 6.4]. In fact, the pairing between $\omega\in\mathbf{CH}^{m,\gamma}([0,1]^{d})$ and $T\in\mathbf{F}_{m}^{\alpha}([0,1]^{d})$ is
@@
More precisely, let $f\colon[0,1]^{d}\to[0,1]^{d^{\prime}}$ be $\alpha$-Hölder continuous. Then $f$ defines a pushforward operator (see [8, Theorem 8.1])

### v2

# arXiv:2602.09151v2[math.FA] 02 Mar 2026

@@ 1. Introduction
The classical Young integral [47] is recovered in the case $d=1$. In recent years, this integral has attracted significant attention. For instance, [1] provides an alternative construction of the two-dimensional Züst integral, while [30] investigates the Hölder regularity of Hölder differential forms arising in the Züst integral. Since Züst’s original work, most contributions have focused on this integral, largely due to its connections with the Gromov–Hölder problem for the Heisenberg group; see, for example, [25, 3].
@@ 2.3. Henstock integral in higher dimension
A natural question is whether one can formulate a divergence theorem for the Henstock integral that is valid for every differentiable vector field. Such a result would constitute a higher-dimensional analog of the fundamental theorem of calculus. Unfortunately, the answer is negative. This non-absolute integral does, however, satisfy a Fubini theorem (see [36, Section 6.1], and a remarkable argument due to W. Pfeffer shows that a gauge integral cannot simultaneously satisfy both a generalized Fubini theorem and divergence theorem, see [39, Example 11.1.2].
@@ 2.4. Charges
The first appearance of charges in the study of non-absolute integrals can be traced back to Mařík [34]. The condition $\sup_{n}\|B_{n}\|<\infty$ in the continuity statement ensures that the $B_{n}$ do not fractalize excessively. It is inspired by continuity properties of the Henstock-Kurzweil integral (see for instance Hake’s theorem [39, Theorem 6.1.6]). Note also that it implies that $\omega(B)=0$ whenever $|B|=0$.
@@ 2.6. Pfeffer integral
$\left|\mathscr{S}(f,\mathscr{P})-g(1)\right|\leq\sum_{i=0}^{n-1}\left|f(x_{i})(a_{i+1}-a_{i})-(g(a_{i+1})-g(a_{i}))\right|\leq\varepsilon.$
@@
The formulation is of Saks-Henstock type because, in general, $\delta$-fine partitions do not always exist. For example, when $d=1$ and $\delta(x)=x/2$ on $[0,1]$, it is impossible to construct a $\delta$-fine partition. However, the well-definedness of this definition is based on a powerful Cousin lemma, due to Howard ([29, Lemma 5] or [41, Lemma 2.6.4]), that guarantees the existence of almost filling $\delta$-fine $\eta$-regular partitions (for $\eta$ smaller than the regularity of cubes), and therefore, the uniqueness of the indefinite integral $\omega$. It is also possible to give a formulation of the Pfeffer integral in terms of Riemann sums, with control over the size of the residual set.
@@
In one dimension, the Pfeffer integral is strictly weaker than the Henstock-Kurzweil integral, essentially because compact intervals are replaced by finite unions of compact intervals, [38, Example 6.9]. This is a drawback, in view of the well-established status of the Henstock-Kurzweil integral. In dimension $d\geq 2$, the Henstock and Pfeffer integrals give rise to different classes of integrable functions, with no inclusion relation between them.
@@ 3.3. A multidimensional Faber-Schauder basis
as in Theorem 3.5. It is possible to make sense of the coefficients $\omega(h_{-1})$ and $\omega(h_{n,k,e})$, because the Haar functions are piecewise constant on dyadic cubes. Moreover, the Hölder control yields estimate of the coefficients $|\omega(h_{n,k,e})|$. Using (3.6), we then show that the series in (3.9) converges in $CH([0,1]^{d})$. Consequently, the left-hand side of (3.9) is well-defined as an element of $CH([0,1]^{d})$ and provides the desired extension. Proving that $\omega$ is actually $\gamma$-fractional requires an interpolation argument that is to be found in [8, Theorem 3]. ∎
@@
Therefore, the space $CH^{\gamma}([0,1]^{d})$ is isomorphic to $\ell^{\infty}$, in the same way that $C^{\gamma}([0,1])$ is related to $\ell^{\infty}$ via the Ciesielski isomorphism. In particular, $CH^{\gamma}([0,1]^{d})$ is a dual Banach space (indeed, a bidual, see [9, Proposition 7.2]).
@@
Finally, we briefly discuss the characterization of $\gamma$-fractional charges. It was shown in [8, Theorem 11] that they are precisely the divergences of $\gamma$-Hölder continuous vector fields. More specifically, the operator $\operatorname{div}$ defined in (3.1) induces a surjective map
@@ 4.1. Whitney’s theory of flat chains and flat cochains
This turns out to be the right choice, as (1) and (2) are easily established. The only difficult step is to prove $\mathbf{F}(T)=0\implies T=0$, which is carried out in [46, Chapter V, §12]. The norm $\mathbf{F}$ is called the flat norm. The following example shows that the flat norm encodes some sort of weak convergence, and why axiom (1) is essential.
@@ 4.3. Fractional currents and fractional charges
Comparing with (4.4), we guess that $\mathbf{CH}^{m,\gamma}([0,1]^{d})$ is the dual of $\mathbf{F}_{m}^{\gamma}([0,1]^{d})$, [9, Theorem 6.4]. In fact, the pairing between $\omega\in\mathbf{CH}^{m,\gamma}([0,1]^{d})$ and $T\in\mathbf{F}_{m}^{\alpha}([0,1]^{d})$ is
@@
More precisely, let $f\colon[0,1]^{d}\to[0,1]^{d^{\prime}}$ be $\alpha$-Hölder continuous. Then $f$ defines a pushforward operator (see [9, Theorem 8.1])